OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i

OpenStudy (anonymous):

9i?

OpenStudy (anonymous):

answer choices: f(x) = x4 - 11x3 + 72x2 - 606x + 10,080 f(x) = x4 - 303x2 + 1212x - 10,080 f(x) = x4 - 11x3 - 72x2 + 606x - 10,080 f(x) = x4 - 58x2 + 1212x - 10,080

OpenStudy (anonymous):

what is this for

OpenStudy (anonymous):

pre calc. its my homework assignment. I tried doing it on Mathway.com but they couldnt even solve it.

OpenStudy (anonymous):

i feel like its impossible

OpenStudy (anonymous):

one min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x^4 -58x^2 + 1212x - 10080

OpenStudy (anonymous):

d

OpenStudy (anonymous):

(x−8)(x+14)(x−(3+9i))(x−(3−9i))

OpenStudy (anonymous):

that was correct. thank you so much (: it helped boost my grade!

OpenStudy (anonymous):

no problem